OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (2, 1) m = 0

OpenStudy (anonymous):

\[y = mx + b.\] \[(2,1) = (x, y)\] \[1 = 0(2) + b.\]

OpenStudy (anonymous):

Now, what do you do after this?

OpenStudy (anonymous):

I have no idea. Im so lost

OpenStudy (anonymous):

What is 2 times 0?

OpenStudy (anonymous):

\[2 \times 0 = ?\]

OpenStudy (anonymous):

What is anything times 0?